# Macro Graph of Experts for Billion-Scale Multi-Task Recommendation

### v1

@@ Abstract.
Graph-based multi-task learning at billion-scale presents a significant challenge, as different tasks correspond to distinct billion-scale graphs. Traditional multi-task learning methods often neglect these graph structures, relying solely on individual user and item embeddings. However, disregarding graph structures overlooks substantial potential for improving performance. In this paper, we introduce the Macro Graph of Expert (MGOE) framework, the first approach capable of leveraging macro graph embeddings to capture task-specific macro features while modeling the correlations between task-specific experts. Specifically, we propose the concept of a Macro Graph Bottom, which, for the first time, enables multi-task learning models to incorporate graph information effectively. We design the Macro Prediction Tower to dynamically integrate macro knowledge across tasks. MGOE has been deployed at scale, powering multi-task learning for the homepage of a leading billion-scale recommender system. Extensive offline experiments conducted on three public benchmark datasets demonstrate its superiority over state-of-the-art multi-task learning methods, establishing MGOE as a breakthrough in multi-task graph-based recommendation. Furthermore, online A/B tests confirm the superiority of MGOE in billion-scale recommender systems.
@@
copyright: noneccs: Information systems Recommender systemsccs: Information systems Business intelligenceccs: Information systems Online advertising
@@ 1. Introduction
Multi-task learning (MTL) lies at the core of nearly all billion-scale recommendation systems (Wang et al., 2023; Caruana, 1997; Kokkinos, 2017), responsible for predicting multiple key tasks such as clicks, likes, and adding to cart (Juan et al., 2016; Zhu et al., 2023). Its goal is to address multiple tasks in real-time to meet diverse user needs. Previous multi-task recommendation models overlook the valuable information embedded within graphs. As shown in Figure 1, their embedding-based multi-task bottoms rely solely on individual user and item embeddings and disregard the rich relational semantics encoded in graph structures, which ultimately results in suboptimal performance. However, incorporating graph structures in MTL presents greater challenges compared to single-task situations due to two primary reasons: 1) online systems have billions of users and items, necessitating efficient and effective handling of this scale, and 2) multiple graph relations are formed by different tasks, requiring appropriate graph management techniques.
The mainstream approach in industrial recommendation typically employs single embeddings for multi-task problems without considering graph structures (Sener and Koltun, 2018; Ruder et al., 2019). ESMM (Ma et al., 2018a) enhances the robustness and improves the overall performance by utilizing the entire exposed sample space for training. MMoE (Ma et al., 2018b) pioneers the use of a basic mixture-of-experts structure, enabling task-specific predictions to learn from other tasks. PLE (Tang et al., 2020) separates task-sharing and task-specific experts to avoid the negative transfer and seesaw phenomenon. STEM (Su et al., 2024) introduces a novel shared and task-specific embeddings paradigm to effectively capture user preferences. Subsequent works focused on designing more sophisticated expert structures, such as multi-layer or task-specialized experts. Other studies explore graph-based multi-task learning; for example, MoGENet (Hu et al., 2022) proposes a multi-channel graph neural networks (GNNs) to model high-order information using user-item bipartite graphs. MMoCEG (Zhang et al., 2023b) introduces graph-based MoE to recognize the commonalities and differences among multiple regions. However, graph-based MTL models often struggle with the high computational burden of GNNs when facing billion-scale graphs.
@@
Figure 1. Comparison of Multi-Task Learning Architectures. (a) Embedding-based MTL relies on feature representations. (b) Graph-based MTL utilizes GNNs to capture relational semantics. (c) Macro Graph MTL introduces macro graph bottom to incorporate macro graph information.
To address these challenges, we introduce Macro Graph of Experts (MGOE), the first graph neural network architecture for billion-scale multi-task recommendation that avoids introducing unbearable computational complexity. We present the Macro Task Merging Graph (MTMG), which extracts multi-task information using a single macro graph. We then describe the design of macro graph experts that learn graph information to predict different tasks. Finally, we introduce the Macro Prediction Tower, which aggregates multi-level information to simultaneously predict multiple tasks.
@@
  • Extensive offline experiments conducted on three public benchmark datasets verify that MGOE outperforms both embedding-based multi-task learning models and graph-based recommendation models. Besides, online A/B test have confirmed the superiority of MGOE in the billion-scale recommender systems.
@@ 2.1. Notations
In this paper, we use $\mathcal{U}$, $\mathcal{I}$, and $\mathcal{T}$ to denote the user set, item set, and task set, respectively. Each task $t\in\mathcal{T}$ is associated with a task-specific interaction matrix ${\mathbf{R}}^{(t)}$. For any user $u\in\mathcal{U}$ and item $i\in\mathcal{I}$, the entry $r^{(t)}_{ui}\in{\mathbf{R}}^{(t)}$ represents the interaction between $u$, $i$, and $t$. Specifically, $r^{(t)}_{ui}=1$ indicates that user $u$ and item $i$ have interacted in task $t$, while $r^{(t)}_{ui}=0$ implies no interaction. For each task $t$, we have a dataset $\mathcal{D}_{t}$, where each element is represented as $(u,i,y_{ui}^{(t)})\in\mathcal{D}_{t}$. Here, the label $y_{ui}^{(t)}\in\{0,1\}$ indicates whether user $u$ interacted (1) or did not interact (0) with item $i$ in task $t$.
@@ 2.2. Multi-task Learning
Multi-task learning (MTL) aims to develop the ability to handle multiple related but distinct tasks simultaneously. In e-commerce scenarios, these tasks may involve predicting whether a user will interact with an item by performing a specific behavior, such as clicking on a product, adding it to a cart, or making a purchase. The objective of MTL is to learn a set of functions $\mathcal{F}$, where each function $f_{t}\in\mathcal{F}$ takes user and item features as input to predict the probability of the corresponding task, which is defined as $f_{t}(\cdot):(u,i)\rightarrow\hat{y}_{ui}^{(t)}$, with $\hat{y}_{ui}^{(t)}\in[0,1]$ representing the predicted probability. Typically, MTL employs the binary cross-entropy loss as the loss function for each task $t$, which is expressed as:
$\mathcal{L}_{t}=-\sum_{(u,i,y_{ui}^{(t)})\in\mathcal{D}_{t}}\left[y_{ui}^{(t)}\log(\hat{y}_{ui}^{(t)})+(1-y_{ui}^{(t)})\log(1-\hat{y}_{ui}^{(t)})\right].$
The overall loss function $\mathcal{L}_{\text{MTL}}$ is then defined as:
(1) $\mathcal{L}_{\text{MTL}}=\sum_{t\in\mathcal{T}}\alpha_{t}\mathcal{L}_{t},$
where $\alpha_{t}$ is the parameter to reflect task $t$’s relative importance.
@@ 2.3. MTL Bottom
The embedding-based MTL bottom serves as the process that transforms feature inputs into fused representations. Typically, for each task, the multi-task bottom takes relevant user micro features ${\bm{e}}_{u}$ and item micro features ${\bm{e}}_{i}$ as input, concatenating them into a vector ${\bm{z}}=[{\bm{e}}_{u};\;{\bm{e}}_{i}]$. This newly formed vector is the input for subsequent expert layers, enabling further feature fusion. Suppose the number of shared experts and task-specific experts is $N_{1}$ and $N_{2}$, respectively. The prediction $\hat{y}_{ui}^{(t)}$ for task $t$ is then expressed as:
(2) $h_{t}\left({\bm{z}}\right)$ $=\sum_{l=1}^{N_{1}}\operatorname{G}_{t,l}^{(1)}\left({\bm{z}}\right)\operatorname{E}_{s,l}\left({\bm{z}}\right)+\sum_{l=1}^{N_{2}}\operatorname{G}_{t,l}^{(2)}\left(\mathbf{z}\right)\operatorname{E}_{t,l}\left({\bm{z}}\right);$
@@
where $\sigma$ denotes the sigmoid function, $\operatorname{E}_{s,l}$ represent the $l$-th shared experts, $\operatorname{E}_{t,l}$ is the $l$-th task-specific experts for task $t$ , $\operatorname{G}_{t,l}^{(1)}(\cdot)$ and $\operatorname{G}_{t,l}^{(2)}(\cdot)$ represent the gating network operations that fuse the outputs from shared experts and task-specific experts, respectively. It is worth noting that embedding-based MTL models separately focus on user and item micro features, neglecting the graph information formed by user-item interactions.
@@
The graph-based bottom typically captures the user-item collaborative connections through graph neural networks (GNNs). Given the input of the $l$-th layer ${\bm{e}}^{l}_{v}$ for a certain user/item node $v$, the output of $(l+1)$-th layer ${\bm{e}}^{l+1}_{v}$ and the prediction $\hat{y}_{ui}^{(t)}$ for task $t$ are represented as:
@@
where ${\mathbf{W}}_{{\mathcal{G}}_{t}}$ is the weight matrix for the task graph ${\mathcal{G}}_{t}$, and $\eta_{v\leftarrow c}$ is implemented using an attention mechanism. Here, $\mathcal{N}(v)$ denotes the set of neighbors of node $v$. However, when dealing with billion-scale task graphs, the graph-based bottom faces significant challenges in computational resources.
@@ 3. Methodology
The overall framework of our proposed MGOE is shown in Figure 2. We begin by introducing the macro task merging graph (MTMG) utilized in this paper. Subsequently, we present our proposed MGOE, which is designed to effectively leverage the macro graph information to enhance recommendation performance.
@@ 3.1. Macro Task Merging Graph
In this section, we focus on constructing MTMG, which provides crucial macro graph information for the subsequent MGOE. The construction of MTMG primarily involves two core steps: first, obtaining a merging interaction matrix, and second, computing the macro nodes and constructing the macro edges, respectively.
@@
In real-world recommendation scenarios, multiple tasks often coexist (e.g., users may click and purchase an item simultaneously). These task combinations represent the co-existence of various interaction patterns between users and items, reflecting different degrees of user preference.
To accurately capture these preferences, we introduce a scoring function to quantify the importance of such task combinations. For any user $u\in\mathcal{U}$ and item $i\in\mathcal{I}$, the possible task combinations set between them are denoted as:
@@
Let $\mathcal{C}=\{\mathbf{C}_{ui}|u\in\mathcal{U},i\in\mathcal{I}\}$ Once we have defined the importance order of each task, we can obtain a rank function for task combination sets (Zhang et al., 2024a), denoted as $\rho(\cdot):\mathcal{C}\to\mathbb{N}^{+}$. We further define a scoring function $g(\cdot):\mathcal{C}\to\mathbb{R}$, which assigns a merging preference score to each task combination, representing its relative importance. Specifically, the merging preference score $s_{ui}$ between user $u$ and item $i$ is defined as:
@@
where $\beta$ is a transformation parameter used to adjust the importance of task combination sets. Then, we arrange these values corresponding to user-item pairs to construct the merging interaction matrix ${\mathbf{S}}$:
@@
where $s_{ui}$ serves as each entry of the matrix, $m=|\mathcal{U}|$ and $n=|\mathcal{I}|$ denotes the number of users and items respectively. The well-designed scoring function enables systematic quantitative evaluations of different task combinations, ensuring that the constructed MTMG effectively captures the complex interactions between users and items. This, in turn, supports the efficient operation of the MGOE. Following the definition of the scoring function, we proceed to construct the nodes and edges within MTMG.
@@
Computing a macro node involves leveraging the merging preference score computed by Equation 6. Existing approaches treat all nodes in the macro graph uniformly, ignoring the rich semantic relationships in different task combinations. To overcome this limitation, we propose Preference-Based Grouping, a novel method that assigns distinct weights to each micro-node based on its merging preference score. Specifically, we construct macro nodes from two complementary perspectives.
Let ${\mathbf{S}}^{\mathcal{U}}={\mathbf{S}}{\mathbf{S}}^{\top}$ and ${\mathbf{S}}^{\mathcal{I}}={\mathbf{S}}^{\top}{\mathbf{S}}$. Then we can obtain macro nodes of by performing clustering on users and items based on ${\mathbf{S}}^{\mathcal{U}}$ and ${\mathbf{S}}^{\mathcal{I}}$, respectively. The following takes clustering of users as an example. Let $d_{vj}$ denote the $(v,j)$-th element of matrix ${\mathbf{S}}^{\mathcal{U}}$. Subsequently, we assign different weights to different micro nodes. We define a merging preference score baseline $s_{\text{base}}$ for node $v$, such that if the sum of the merging preference score values exceeding $s_{\text{base}}$ surpasses a preset threshold $\delta_{1}$, the node is considered active, and a relatively high weight $w_{\text{high}}$ is assigned. Conversely, if the sum falls below a certain value $\delta_{2}\,\left(\delta_{2}\leq\delta_{1}\right)$, the node is regarded as inactive, and a relatively low weight $w_{\text{low}}$ is assigned. For all other cases, a weight of 1 is assigned. Mathematically, the weight assigned to micro nodes $v$ is defined as:
(8) $w_{v}=\begin{cases}w_{\text{high}}\,,&\sum_{j=1}^{n}\mathbbm{1}_{\{d_{vj}\geq s_{\text{base}}\}}\geq\delta_{1}\\ w_{\text{low}}\,,&\sum_{j=1}^{n}\mathbbm{1}_{\{d_{vj}\geq s_{\text{base}}\}}\leq\delta_{2}\\ 1,&\text{otherwise}\end{cases}\,,$
@@
where $\bm{\mu}_{k}$ is the centroid of cluster $k$, $\mathcal{V}_{k}$ denotes the set of micro-nodes assigned to cluster $k$, $w_{v}$ is the weight associated with node $v$, and $\mathbf{d}_{v}=(d_{v1},d_{v2}\cdots,d_{vn})^{\top}$ is the $v$-th row of ${\mathbf{S}}^{\mathcal{U}}$. The clustering of items can be performed using the same method based on ${\mathbf{S}}^{\mathcal{I}}$. Preference-based grouping ensures that micro-nodes with larger weights exert greater influence on the centroid’s position, thereby making the centroid determination more flexible and better aligned with users’ actual interests and task patterns.
@@
Our proposed MTMG can be formally defined as $\widetilde{{\mathcal{G}}}_{\text{MTMG}}=\left(\widetilde{\mathcal{U}},\widetilde{\mathcal{I}},\widetilde{{\mathbf{S}}}\right)$, where $\widetilde{\mathcal{U}}$ represents the set of macro user nodes and $\widetilde{\mathcal{I}}$ denotes set of the macro item nodes. Each macro node is associated with a trainable macro embedding $\widetilde{\textbf{E}}$, and its macro neighbors are represented as $\widetilde{\mathcal{N}}$. MTMG captures task patterns from a macro perspective and provides macro graph information for the subsequent macro graph bottom.

### v4

@@ Abstract.
Graph-based multi-task learning at billion-scale presents a significant challenge, as different tasks correspond to distinct billion-scale graphs. Traditional multi-task learning methods often neglect these graph structures, relying solely on individual user and item embeddings. However, disregarding graph structures overlooks substantial potential for improving performance. In this paper, we introduce the Macro Graph of Experts (MGOE) framework, the first approach capable of leveraging macro graph embeddings to capture task-specific macro features while modeling the correlations between task-specific experts. Specifically, we propose the concept of a Macro Graph Bottom, which, for the first time, enables multi-task learning models to incorporate graph information effectively. We design the Macro Prediction Tower to dynamically integrate macro knowledge across tasks. MGOE has been deployed at scale, powering multi-task learning for a leading billion-scale recommender system, Alibaba. Extensive offline experiments conducted on three public benchmark datasets demonstrate its superiority over state-of-the-art multi-task learning methods, establishing MGOE as a breakthrough in multi-task graph-based recommendation. Furthermore, online A/B tests confirm the superiority of MGOE in billion-scale recommender systems.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Republic of Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770854.3783958isbn: 979-8-4007-2258-5/2026/08ccs: Information systems Recommender systemsccs: Information systems Business intelligenceccs: Information systems Online advertising
@@ 1. Introduction
Multi-task learning (MTL) lies at the core of nearly all billion-scale recommendation systems (Wang et al., 2023; Caruana, 1997; Kokkinos, 2017; Zhang et al., 2023b), responsible for predicting multiple key tasks such as clicks, likes, and adding to cart (Juan et al., 2016; Zhu et al., 2023). Its goal is to address multiple tasks in real-time to meet diverse user needs. Previous multi-task recommendation models overlook the valuable information embedded within graphs. As shown in Figure 1, their embedding-based multi-task bottoms rely solely on individual user and item embeddings and disregard the rich relational semantics encoded in graph structures, which ultimately results in suboptimal performance. However, incorporating graph structures in MTL presents greater challenges compared to single-task situations due to two primary reasons: 1) online systems have billions of users and items, necessitating efficient and effective handling of this scale, and 2) multiple graph relations are formed by different tasks, requiring appropriate graph management techniques.
The mainstream approach in industrial recommendation typically employs single embeddings for multi-task problems without considering graph structures (Sener and Koltun, 2018; Ruder et al., 2019; Shen et al., 2025). ESMM (Ma et al., 2018b) enhances the robustness and improves the overall performance by utilizing the entire exposed sample space for training. MMoE (Ma et al., 2018a) pioneers the use of a basic mixture-of-experts structure, enabling task-specific predictions to learn from other tasks. PLE (Tang et al., 2020) separates task-sharing and task-specific experts to avoid the negative transfer and seesaw phenomenon. STEM (Su et al., 2024) introduces a novel shared and task-specific embeddings paradigm to effectively capture user preferences. Subsequent works focused on designing more sophisticated expert structures, such as multi-layer or task-specialized experts. Other studies explore graph-based multi-task learning; for example, MoGENet (Hu et al., 2022) proposes a multi-channel graph neural networks (GNNs) to model high-order information using user-item bipartite graphs. MMoCEG (Zhang et al., 2023a) introduces graph-based MoE to recognize the commonalities and differences among multiple regions. However, graph-based MTL models often struggle with the high computational burden of GNNs when facing billion-scale graphs.
Figure 1. Comparison of Multi-Task Learning Architectures, where ”G” denotes a gating mechanism and ”MGE” represents the proposed Macro Graph Expert. (a) Embedding-based MTL relies on feature representations. (b) Graph-based MTL utilizes GNNs to capture relational semantics. (c) Macro Graph MTL introduces macro graph bottom to incorporate macro graph information.
@@
By addressing these challenges, we introduce Macro Graph of Experts (MGOE), the first graph neural network architecture for billion-scale multi-task recommendation that avoids introducing unbearable computational complexity. We present the Macro Task Merging Graph (MTMG), which extracts multi-task information using a single macro graph. We then describe the design of macro graph experts that learn graph information to predict different tasks. Finally, we introduce the Macro Prediction Tower, which aggregates multi-level information to simultaneously predict multiple tasks.
@@
  • Extensive offline experiments conducted on three public benchmark datasets verify that MGOE outperforms both embedding-based multi-task learning models and graph-based recommendation models.  [^1] Source code is available at https://github.com/RainmannnnN/MGOE Besides, online A/B test have confirmed the superiority of MGOE in the billion-scale recommender systems.
@@ 2.1. Notations
We use $\mathcal{U}$, $\mathcal{I}$, and $\mathcal{T}$ to denote the user set, item set, and task set, respectively. Each task $t\in\mathcal{T}$ is associated with a task-specific interaction matrix ${\mathbf{R}}^{(t)}$. For any user $u\in\mathcal{U}$ and item $i\in\mathcal{I}$, the entry $r^{(t)}_{ui}\in{\mathbf{R}}^{(t)}$ represents the interaction between $u$ and $i$ under task $t$. Specifically, $r^{(t)}_{ui}=1$ indicates that user $u$ has interacted with item $i$, while $r^{(t)}_{ui}=0$ indicates no interaction. Each task $t$ is also associated with a dataset $\mathcal{D}_{t}$, where each element is represented as $(u,i,y_{ui}^{(t)})\in\mathcal{D}_{t}$. Here, the label $y_{ui}^{(t)}\in\{0,1\}$ indicates whether user $u$ interacted (1) or did not interact (0) with item $i$ in task $t$. $r^{(t)}_{ui}$ and $y_{ui}^{(t)}$ are numerical equivalent but have representations.
@@ 2.2. Multi-Task Learning
Multi-task learning aims to develop the ability to handle multiple related but distinct recommendation tasks simultaneously, which may involve predicting whether a user will interact with an item by performing a specific behavior (e.g., clicking on a product, adding it to a cart, or making a purchase). While these tasks share foundational latent factors, including user preference structures and item attribute representations, each corresponds to distinct decision stages and motivations, exhibiting significant task-specific characteristics. The target of MTL is to learn a functions set $\mathcal{F}$, where each function $f_{t}\in\mathcal{F}$ takes user and item features as input to predict the probability $\hat{y}_{ui}^{(t)}\in[0,1]$ of potential interaction under the corresponding task, which is defined as $f_{t}(\cdot):(u,i)\rightarrow\hat{y}_{ui}^{(t)}$. Typically, MTL employs the binary cross-entropy loss as the loss function for each task $t$:
(1) $\mathcal{L}_{t}=-\sum_{(u,i,y_{ui}^{(t)})\in\mathcal{D}_{t}}\left[y_{ui}^{(t)}\cdot\log(\hat{y}_{ui}^{(t)})+(1-y_{ui}^{(t)})\cdot\log(1-\hat{y}_{ui}^{(t)})\right].$
The overall loss function $\mathcal{L}_{\text{MTL}}$ is obtained by weighted sum:
(2) $\mathcal{L}_{\text{MTL}}=\sum_{t\in\mathcal{T}}\alpha_{t}\cdot\mathcal{L}_{t},$
where parameter $\alpha_{t}$ denotes the relative importance of task $t$.
@@ 2.3. MTL Bottom
The embedding-based MTL bottom transforms input features into fused representations. For each task, the multi-task bottom takes vector ${\bm{z}}=[{\bm{e}}_{u};\;{\bm{e}}_{i}]$ concatenated by relevant micro features from user ${\bm{e}}_{u}$ and item ${\bm{e}}_{i}$ as input, enabling further feature fusion for subsequent expert layers. Suppose the number of shared experts and task-specific experts is $N_{1}$ and $N_{2}$, the predicted probability $\hat{y}_{ui}^{(t)}$ for task $t$ can be expressed by:
(3) $h_{t}\left({\bm{z}}\right)$ $=\sum_{l=1}^{N_{1}}\operatorname{G}_{t,l}^{(1)}\left({\bm{z}}\right)\cdot\operatorname{E}_{s,l}\left({\bm{z}}\right)+\sum_{l=1}^{N_{2}}\operatorname{G}_{t,l}^{(2)}\left({\bm{z}}\right)\cdot\operatorname{E}_{t,l}\left({\bm{z}}\right);$
@@
where $\sigma$ denotes the sigmoid function, $\operatorname{E}_{s,l}$ and $\operatorname{E}_{t,l}$ represent the $l$-th shared experts and task-specific experts for task $t$, respectively. $\operatorname{G}_{t,l}^{(1)}(\cdot)$ and $\operatorname{G}_{t,l}^{(2)}(\cdot)$ denote the gating network operations that fuse the outputs from the corresponding experts. Note that the embedding-based MTL bottom focuses on micro features separately, neglecting the information induced by user-item interactions.
@@
The graph-based bottom captures the user-item collaborative connections through GNNs. Given the input of the $l$-th layer ${\bm{e}}^{l}_{v}$ for a certain user/item node $v$, the output of $(l+1)$-th layer ${\bm{e}}^{l+1}_{v}$ and the prediction $\hat{y}_{ui}^{(t)}$ for task $t$ are:
@@
where ${\mathbf{W}}_{{\mathcal{G}}_{t}}$ is the weight matrix of the task graph ${\mathcal{G}}_{t}$, $\eta_{v\leftarrow c}$ is implemented via attention, and $\mathcal{N}(v)$ denotes the set of neighbors of node $v$. When applied to billion-scale task graphs, the graph-based bottom encounters significant computational challenges.
2.4. Micro Graph
A micro graph is a fine-grained structure that models user-item interactions individually, where users and items are treated as micro nodes. Micro edges are instantiated through actual interactions (e.g., purchases) and are represented by an interaction matrix. Micro features characterize both micro nodes and edges. The trainable embeddings of micro nodes encode latent semantic information, while the entries of the interaction matrix $\mathbf{R}$} quantify the properties of micro edges. This structure captures granular user-item relationships and serves as a foundational component in GNN-based recommendation models, enabling fine-grained interaction analysis and prediction. Formally, a micro graph is denoted as $\mathcal{G}=(\mathcal{U},\mathcal{I},{\mathbf{R}})$.
@@ 3. Methodology
The overall framework of our proposed MGOE is illustrated in Figure 2. In this section, we firstly introduce the utilized MTMG. Subsequently, we present the MGOE with detailed structure.
@@ 3.1. Macro Task Merging Graph
We begin by constructing MTMG, which provides essential macro-level graph information for the subsequent MGOE modeling. The construction of MTMG primarily involves two core steps: (1) obtaining a merged interaction matrix, and (2) computing the macro nodes and constructing the macro edges in order.
@@
In real-world scenarios, multiple tasks often coexist (e.g., users may click on and purchase an item simultaneously). These task combinations reflect the coexistence of diverse interaction patterns between users and items, indicating varying degrees of user preference. To accurately capture these preferences, we introduce a scoring function to quantify the importance level of such task combinations. For any user $u\in\mathcal{U}$ and item $i\in\mathcal{I}$, the potential mutual task combinations set is represented by:
@@
Let the collection of $\mathbf{C}_{ui}$ as $\mathcal{C}=\{\mathbf{C}_{ui}\mid\ u\in\mathcal{U},i\in\mathcal{I}\}$. Once we have defined the importance order of each task, we can obtain a rank function for task combination sets (Zhang et al., 2024), denoted as $\rho(\cdot):\mathcal{C}\to\mathbb{N}^{+}$. We further define a scoring function $g(\cdot):\mathcal{C}\to\mathbb{R}$, which assigns a merging preference score to each task combination, representing its relative importance. Specifically, the merging preference score $s_{ui}$ between user $u$ and item $i$ is defined as:
@@
where $\beta$ is a transformation parameter used to adjust the importance of $\mathbf{C}_{ui}$. Then, we arrange these values corresponding to user-item pairs to construct the merging interaction matrix $\mathbf{S}$}:
@@
where $s_{ui}$ denotes each entry of the matrix, and $m=\left|\mathcal{U}\right|$ and $n=\left|\mathcal{I}\right|$ represent the number of users and items, respectively. The scoring function enables systematic and quantitative evaluation of task combinations, ensuring that MTMG accurately captures complex patterns in user-item interactions. This enhances both the expressiveness of the macro graph and the efficiency of MGOE in downstream MTL. Following the definition of the scoring function, we proceed to construct the nodes and edges within MTMG.
@@
Computing a macro node involves leveraging the merging preference score computed by Equation 7. Existing approaches treat all nodes in the macro graph uniformly, ignoring the rich semantic relationships in different task combinations. To overcome this limitation, we propose Preference-Based Grouping, a novel method that assigns distinct weights to each micro node based on its merging preference score. Specifically, we construct macro nodes from two complementary perspectives.
Let ${\mathbf{S}}^{\mathcal{U}}={\mathbf{S}}{\mathbf{S}}^{\top}$ and ${\mathbf{S}}^{\mathcal{I}}={\mathbf{S}}^{\top}{\mathbf{S}}$. Then we can obtain macro nodes of by performing clustering on users and items based on ${\mathbf{S}}^{\mathcal{U}}$ and $\mathbf{S}$$^{\mathcal{I}}$, respectively. The following takes clustering of users as an example. Let $d_{vj}$ denote the $(v,j)$-th element of matrix ${\mathbf{S}}^{\mathcal{U}}$. Subsequently, we assign different weights to different micro nodes. We define a merging preference score baseline $s_{0}$ for node $v$, such that if the sum of the merging preference score values exceeding $s_{0}$ surpasses a preset threshold $\delta_{1}$, the node is considered active, and a relatively high weight $w^{+}$ is assigned. Conversely, if the sum falls below a certain value $\delta_{2}\,\left(\delta_{2}\leq\delta_{1}\right)$, the node is regarded as inactive, and a relatively low weight $w^{-}$ is assigned. For all other cases, a weight of 1 is assigned. Mathematically, the weight assigned to micro nodes $v$ is defined as:
(9) $w_{v}=\begin{cases}w^{+}\,,&\sum_{j=1}^{m}\mathbbm{1}_{\{d_{vj}\geq s_{0}\}}\geq\delta_{1}\\ w^{-}\,,&\sum_{j=1}^{m}\mathbbm{1}_{\{d_{vj}\geq s_{0}\}}\leq\delta_{2}\\ 1,&\text{otherwise}\end{cases}\,$,
@@
where $\bm{\mu}_{k}$ is the centroid of cluster $k$, $\mathcal{V}_{k}$ denotes the set of micro nodes assigned to cluster $k$, $w_{v}$ is the weight associated with node $v$, and $\mathbf{d}_{v}=(d_{v1},d_{v2}\cdots,d_{vm})^{\top}$ is the $v$-th row of ${\mathbf{S}}^{\mathcal{U}}$. The clustering of items can be performed using the same method based on $\mathbf{S}$$^{\mathcal{I}}$. Preference-based grouping ensures that micro nodes with larger weights exert greater influence on the centroid’s position, thereby making the centroid determination more flexible and better aligned with users’ actual interests and task patterns.
@@
[…]